14. A box is raised using a rope. If the acceleration of the box, is 6.1 m/s2 (upward) and the tension in the rope is 61 N, what is the mass of the box (in kg)?14. A box is raised using a rope. If the accelerat
Answer #1

Here accelaration due to gravity and accelaration of rope are in opposite direction

\Rightarrow Resultant acc = aR = (g+a)

tension in rope ( T ) = Weight of object ( W )

T = mass of object * Resultant accelaration

T = m * ( g + a)

61 = m * ( 9.8 + 6.1 )

\Rightarrow Mass of box = m = 61 / 15.9 = 3.83 kg